# Absence of charged pion condensation in a magnetic field with parallel rotation

### v2

@@ I Introduction
This paper is organized as follows. In Sec. II we set up a field theoretical model for relativistic bosons in PRM and evaluate the partition function in the imaginary-time formalism. In Sec. III we study the critical temperature for the non-interacting case. In Sec. IV we study the effect of phase fluctuation on the order parameter and the off-diagonal long-range order. We summarize in Sec. V. The natural units $c=\hbar=k_{\rm B}=1$ is used throughout.
@@ II The partition function at finite temperature
where $\beta=1/T$, $\hat{N}$ is the conserved charge associated with the U(1)(1) symmetry,
@@ III Non-Interacting Bosons in PRM
For non-interacting many-boson systems, the simple way to judge whether BEC exists is to study the conserved quantities of the system. Without loss of generality, we consider the case $\mu>0$. It is clear that when $\mu=m$, the Bose-Einstein distribution function with energy $E_{\bf k}-\mu$ develops a singularity at ${\bf k}=0$. However, the critical temperature $T_{\rm c}$ cannot be determined as a function of $\mu$, since $\mu$ is fixed to be $\mu=m$ in the BEC phase and hence no longer a thermodynamic variable. The correct way to determine $T_{\rm c}$ in this non-interacting system is to impose the conserved quantities. The U(1)(1) charge is actually composed from two contributions,
@@
where $Q_{\rm c}$ is from the condensed particles and the other term is from thermally excited particles and antiparticles. At $T=T_{\rm c}$, we have $Q_{\rm c}=0$ and $\mu=m$. Thus the critical temperature is determined by
@@
In the nonrelativistic limit, $Q$ reduces to the particle number and we recover the well-known result in the textbook. We note that, the existence of a nonzero $T_{\rm c}$ is due to the fact that the integral (29) is safe in the infrared,
@@
To determine the radial eigenfunction $\varphi(\rho)\equiv\varphi_{sl}(\rho)$, an boundary condition is needed. There are two kinds of boundary conditions that are usually used: One is to impose the Dirichlet boundary condition at $\rho=R$, $\varphi_{sl}(R)=0$; the other is to require that the radial part $\varphi_{sl}(\rho)$ is square-integrable over the entire regime $0\leq\rho<\infty$. If we choose the first one, the radial solution takes the form
@@
where $L_{s}^{(|l|)}$ are the Laguerrel polynomials ($s=0,1,2,\cdots$). The corresponding eigenvalues (Landau levels) take a very simple analytical form
@@
Here $L_{z\rm c}$ denotes the angular momentum from the condensed particles. At $T=T_{\rm c}$, we have $L_{z\rm c}=0$ and $\Omega=\Omega_{\rm c}$. Thus the critical temperature $T_{\rm c}$ is determined by
@@
It is infrared divergent. Therefore, $T_{\rm c}$ must vanish. In Fig. 1, we calculate the angular speed $\Omega$ as a function of the temperature $T$ for different values of $L_{z}$. The numerical results indicate that for arbitrary values of $L_{z}$, angular speed $\Omega$ approaches its critical value $\Omega_{\rm c}$ at $T=0$. Thus the system does not Bose condense at any nonzero temperature $T\neq 0$. The physical picture now becomes clear: In a magnetic field, the motion of the particles becomes effectively one-dimensional.
@@
Therefore, we have $L_{z\rm c}=NQ_{\rm c}$. At $T=T_{\rm c}$, $v=0$. Thus we obtain the same conclusion $T_{\rm c}=0$.
In the above calculations we have used a convenient boundary condition which makes the results analytical. Obviously, other boundary conditions (such as the Dirichlet boundary condition) do not lead to a qualitatively different conclusion. The reason for the absence of BEC is the quasi-one-dimensional feature of the system, which is independent of the boundary condition.
IV Interacting Bosons in RPM
@@
At $T\neq 0$, it vanishes exponentially as $|z_{1}-z_{2}|\to\infty$, indicating the absence of ODLRO at any nonzero temperature. This result is in accordance to the Coleman-Mermin-Wagner-Hohenberg theorem CMWH : The Goldstone boson is quasi-one-dimensional and hence the formation of symmetry-breaking homogeneous long-range order along the $z$ direction is forbidden.
@@ V Summary
Charged pion condensation driven by rotation in a parallel magnetic field was previously proposed by Liu and Zahed PRL120:032001 . In this work, we have studied the critical temperature of the charged pion condensation and have shown that the charged pion condensation is actually absent. The charged pions are described by a complex scalar field with a quartic self-interaction. We evaluate the partition function and the critical temperature by using the standard field theoretical method. For non-interacting bosons, the critical temperature can only be determined for a system with fixed angular momentum. We find that the critical temperature of the non-interacting system is zero due to the fact that the system is quasi-one-dimensional, indicating that non-interacting bosons do not condense. For interacting bosons, we show that the order parameter is zero and the off-diagonal long-range order is absent at any nonzero temperature because of the quasi-one-dimensional feature, in accordance with the Coleman-Mermin-Wagner-Hohenberg theorem.
Acknowledgments: The work is supported by the National Natural Science Foundation of China (Grant Nos. 11775123).
@@ Appendix A Analysis of the effective potential
The phase transition is second-order phase since $C_{N,N,N,N}>0$. For instance, setting
@@ Appendix B Calculation of the integral in (100)
In (100) we need to evaluate the integral

### v3

@@ I Introduction
This paper is organized as follows. In Sec. II we set up a field theoretical model for relativistic bosons in PRM and evaluate the partition function in the imaginary-time formalism. In Sec. III we study the critical temperature for the non-interacting case. In Sec. IV we study the effect of phase fluctuation on the order parameter and the off-diagonal long-range order. We summarize in Sec. V. The natural units $c=\hbar=k_{\rm B}=1$ are used throughout.
@@ II The partition function at finite temperature
where $\beta=1/T$, $\hat{Q}$ is the conserved charge associated with the U(1)(1) symmetry,
@@ III Non-Interacting Bosons in PRM
For non-interacting many-boson systems, the simple way to judge whether BEC exists is to study the conserved quantities of the system. Without loss of generality, we consider the case $\mu>0$. It is clear that when $\mu=m$, the Bose-Einstein distribution function with energy $E_{\bf k}-\mu$ develops a singularity at ${\bf k}=0$. However, the critical temperature $T_{\rm c}$ cannot be determined as a function of $\mu$, since $\mu$ is fixed to be $\mu=m$ in the BEC phase and hence no longer a thermodynamic variable. The correct way to determine $T_{\rm c}$ in this non-interacting system is to impose the conserved quantities. The U(1)(1) charge is actually composed of two contributions,
@@
where $Q_{\rm c}$ is from the condensed particles, and the other term is from thermally excited particles and antiparticles. At $T=T_{\rm c}$, we have $Q_{\rm c}=0$ and $\mu=m$. Thus, the critical temperature is determined by
@@
In the nonrelativistic limit, $Q$ reduces to the particle number, and we recover the well-known result in the textbook. We note that, the existence of a nonzero $T_{\rm c}$ is due to the fact that the integral (29) is safe in the infrared,
@@
To determine the radial eigenfunction $\varphi(\rho)\equiv\varphi_{sl}(\rho)$, an boundary condition is needed. Two kinds of boundary conditions are usually used: one is to impose the Dirichlet boundary condition at $\rho=R$, $\varphi_{sl}(R)=0$; the other is to require that the radial part $\varphi_{sl}(\rho)$ is square-integrable over the entire regime $0\leq\rho<\infty$. If we choose the first one, the radial solution takes the form
@@
where $L_{s}^{(|l|)}$ are the Laguerre polynomials ($s=0,1,2,\cdots$). The corresponding eigenvalues (Landau levels) take a very simple analytical form
@@
Here $L_{z\rm c}$ denotes the angular momentum from the condensed particles. At $T=T_{\rm c}$, we have $L_{z\rm c}=0$ and $\Omega=\Omega_{\rm c}$. Thus, the critical temperature $T_{\rm c}$ is determined by
@@
It is infrared divergent. Therefore, $T_{\rm c}$ must vanish. In Fig. 1, we calculate the angular speed $\Omega$ as a function of the temperature $T$ for different values of $L_{z}$. The numerical results indicate that for arbitrary values of $L_{z}$, angular speed $\Omega$ approaches its critical value $\Omega_{\rm c}$ at $T=0$. Thus, the system does not Bose condense at any nonzero temperature $T\neq 0$. The physical picture now becomes clear: In a magnetic field, the motion of the particles becomes effectively one-dimensional.
@@
Therefore, we have $L_{z\rm c}=NQ_{\rm c}$. At $T=T_{\rm c}$, $v=0$. Thus, we obtain the same conclusion $T_{\rm c}=0$.
In the above calculations, we have used a convenient boundary condition that makes the results analytical. Obviously, other boundary conditions (such as the Dirichlet boundary condition) do not lead to a qualitatively different conclusion. The reason for the absence of BEC is the quasi-one-dimensional feature of the system, which is independent of the boundary condition.
IV Interacting Bosons in PRM
@@
At $T\neq 0$, it vanishes exponentially as $|z_{1}-z_{2}|\to\infty$, indicating the absence of ODLRO at any nonzero temperature. This result is in accordance to the Coleman-Mermin-Wagner-Hohenberg theorem CMWH : The Goldstone boson is quasi-one-dimensional, and hence the formation of symmetry-breaking homogeneous long-range order along the $z$ direction is forbidden.
Beyond the discussions above, one possible concern is that, in principle, the $U(1)$ gauge field should also be treated as a dynamical degree of freedom. In this case, the gauge field could absorb the would-be Goldstone mode and become massive via the Anderson-Higgs mechanism. Since there is no physical Goldstone mode, our above conclusion appears invalid. However, this plausible concern does not influence the result. We emphasize here that the ground state $\Phi_{c}$ is inhomogeneous in the x-y plane, and hence there is no translational symmetry in the x-y plane. For this reason, the gauge field propagator is also quasi-one-dimensional. Qualitatively, the long-range behavior of the two-point correlation function of such a massive quasi-one-dimensional gauge field is
$\lim_{|z_{1}-z_{2}|\to\infty}\langle A_{3}(z_{1})A_{3}(z_{2})\rangle\sim Ke^{-M|z_{1}-z_{2}|},$ (102)
where $A_{3}$ is the z-component of the gauge field, $K$ is a constant added for dimensional consistency, and $M$ is the mass acquired through the Anderson-Higgs mechanism. The ODLRO between two points that differ only in their z-coordinate could be calculated similarly:
$\langle\Phi^{*}(X_{1})\Phi(X_{2}){\rm e}^{iq\int_{X_{1}}^{X_{2}}A_{\mu}dx^{\mu}}\rangle\propto\exp\Bigl[-\frac{1}{2}\int_{z_{1}}^{z_{2}}\mathrm{d}z\int_{z_{1}}^{z_{2}}\mathrm{d}z^{\prime}\langle A_{3}(z)A_{3}(z^{\prime})\rangle\Bigr]\propto\exp\Bigl[-\frac{K|z_{1}-z_{2}|}{2M}\Bigr].$ (103)
Then the result in Eq. (101) is qualitatively reproduced, which agrees with our expectation that the ODLRO in Eq. (96) is gauge-invariant. However, owing to the use of a rotating coordinate frame, the quantitative analytical computation of the gauge field correlation functions proves exceedingly challenging. Therefore, we keep the phase fluctuation as a dynamical field and integrate over it when calculating the ODLRO, similar to the standard paradigm for studying superconductivity in condensed matter theory.
@@ V Summary
Charged pion condensation driven by rotation in a parallel magnetic field was previously proposed by Liu and Zahed PRL120:032001 . In this work, we have studied the critical temperature of charged pion condensation and have shown that it is absent. The charged pions are described by a complex scalar field with a quartic self-interaction. We evaluate the partition function and the critical temperature using the standard field-theoretical method. For non-interacting bosons, the critical temperature can only be determined for a system with fixed angular momentum. We find that the critical temperature of the non-interacting system is zero due to the fact that the system is quasi-one-dimensional, indicating that non-interacting bosons do not condense. For interacting bosons, we show that the order parameter is zero and the off-diagonal long-range order is absent at any nonzero temperature because of the quasi-one-dimensional feature, in accordance with the Coleman-Mermin-Wagner-Hohenberg theorem.
Acknowledgments: The work is supported by the National Natural Science Foundation of China (Grant No. 11775123).
@@ Appendix A Analysis of the effective potential
The phase transition is second-order since $C_{N,N,N,N}>0$. For instance, setting
@@
A more recent study PhysRevD.109.056024 indicates that in some parameter regions, the system might have more than one $v_{l}$ is non-negligible, and such a ground state is called "vortex lattice". An exact analytical treatment of this situation is exceedingly difficult. Nevertheless, the following argument can be made: the vortex lattice breaks the translational symmetry of the ground state more strongly than a single vortex, so the associated Goldstone mode and gauge field propagator remain quasi-one-dimensional. Therefore, our conclusion remains valid.
@@ Appendix B Calculation of the integral in (100)
In (100), we need to evaluate the integral